In glass prism, when the angle of incident increases, the angle of refraction ___

1
Expert's answer
2022-06-28T08:19:12-0400

The Snell's law says:

"n_i\sin\theta_i=n_r\sin\theta_r""\theta_r=\sin^{-1}(n_i\sin\theta_i/n_r)"

Therefore, when the angle of incident increases, the angle of refraction increases too.